LEP EXPERIMENTS COMMITTEE

Minutes of the 38th meeting,
16 September 1994






OPEN SESSION

  1. LEP machine status

R. Bailey reported on LEP operations so far this year. Peak luminosities are routinely above 1.5 x 1031 cm-2s-1 and a new record of 2.2 x 1031 cm-2s-1 has been achieved. The machine is currently operating at an integrated luminosity of 600 nb-1 per day, and 48 pb-1 has been accumulated in 1994 up to 12 September. Although some incremental improvements could be envisaged, this performance essentially represents maturity for the 8-bunch pretzel optics (typical tune shifts are 0.035, with a best of 0.043). Each of the LEP detectors has encountered some background problems but they have been solved with the exception of L3, which still suffers from frequent background spikes and Time Expension Chambertrips.
 

2. Report from the bunch-trains working group

T. Camporesi presented the report from the bunch train working group (LEPC 94-13/M110). The design train lengths are 425 ns (4 x 2 bunches per beam), 670 ns (4 x 3) and 741 ns (4 x 4). The four LEP detectors have indicated they will be prepared to handle all of these trains by the beginning of the 1995 run, with the exception of ALEPH which cannot accommodate at present a train length greater than 680 ns (i.e. the 4th bunch in the 4 x 4 scheme would not contribute). Active work is going on in the collaboration to try to find a solution to this problem. So far, all machine tests of bunch-train operation have been positive -in particular an improved asymmetric beam separation near the interaction region has provided a strong reduction of detector backgrounds. The definitive machine tests with a train of 1 x n (n = 2, 3 or 4) bunches providing collisions in ALEPH and DELPHI will take place after the October technical stop.
 

3. Synchrotron radiation masks for LEP 2

G. Von Holtey presented a status report on the preparations for the additional radiation masks required for LEP 2 operation. Highly successful tests have been made in DELPHI of the new synchrotron radiation masks and shielding at z = 2.3 m. Photon background rates in DELPHI have been reduced by up to an order of magnitude, without any increase in the off-momentum particle background. The masks produce no deterioration in LEP machine performance. The designs of synchrotron radiation masks for the other three LEP detectors are well advanced; the masks for OPAL will be installed in the 94/95 shutdown and for ALEPH and L3 in the 95/96 shutdown.
 

4. LEP 2 physics workshop update

F. Zwirner outlined the plans for the workshop on physics at LEP 2. The organising committee (chaired by G. Altarelli) is planning three, or perhaps four, general meetings from January to October 1995 leading to the publication of a CERN yellow report by the end of the year. The individual working groups are organised into four main groups: standard physics, new physics, event generators and physics-machine interface. Convenors and contact persons for each of these groups are being finalised over the next few weeks.
 

5. Presentations of experimental proposals

A. Blondel presented a proposal to test longitudinal spin polarization at LEP (LEPC 94-8/ P8).

A. Ball presented a proposal to install scintillating tiles in the OPAL endcap (LEPC 94-9/P3).

H. Wachsmuth presented a proposal to search for cosmic ray coincidences with detectors in LEP; CosmoLEP (LEPC 94-10/M109).
 

CLOSED SESSION

Present: J. Allaby, W. Bartel, M. Calvetti, T. Camporesi, G. Coignet, J.D. Dowell (Chairman), L. Foà, B. Foster, P. Hansen, W. Hollik, K. Hübner, P.G. Innocenti, J. Kirkby (Secretary), K.H. Kissler, M. Klein, B. Mansoulié, G. Matthiae, D.J. Miller, S. Myers, J. Panman, D. Schlatter, A. Weidberg, H. Wenninger, D.O. Williams and F. Zwirner.

Apologies: G. Kantardjian.
 

1. Approval of the minutes of the 37th meeting

The minutes of the 37th meeting (LEPC 94-11, LEPC 37) were approved without modifications.
 

2. Chairman's report

The Chairman welcomed the new members of the committee: M. Calvetti, P. Hansen, M. Klein, B. Mansoulié and J. Panman.

The Chairman thanked the RF group for the committeeís visit of the LEP 2 superconducting cavities test and assembly area on 15 September. The visit had provided the committee with a deeper understanding and appreciation of the work underway to prepare the LEP 2 superconducting cavities.

The Chairman joined with the committee in congratulating the LEP team for the record performance of the machine. It appears very likely that the target for 1994 of 60 pb-1 will be exceeded.

He also reported that the Research Board had approved the recommendation concerning bunch trains made at the last LEPC meeting.
 

3. Discussion on bunch-trains

The committee took note of the encouraging machine tests related to the operation of LEP with bunch trains, and looks forward to hearing the results of the definitive tests of up to 1 x 4 bunches at the next meeting in December.
 

4. Discussion of the LEP 2 detector masks

The committee was impressed by the excellent results obtained with the new LEP 2 synchrotron radiation masks in DELPHI and congratulated the machine and experimental teams responsible. Sufficient information now exists that the committee encouraged all LEP detectors to proceed with their plans to install synchrotron radiation masks and shielding near z = 2.3 m.
 

5. Discussion of the Proposal LEPC 94-8/P8: A Test of Longitudinal Spin Polarization at LEP

The three referees (W. Bartel, W. Hollik and G. Matthiae) presented their report on this proposal and thanked J. Jowett for providing expert advice on the machine-related aspects. Recent measurements have shown a substantial reduction of transverse polarization at the Z0 peak due to collisions (beam-beam interaction): with the present optics, the polarization is reduced from 20%, before colliding, to an asymptotic value of 5% after colliding. In view of the importance of first observing high (~50%) transverse beam polarization at the Z0 with high luminosity, the committee recommended a two-day machine test be taken in 1994 (from physics running time) at an energy of peak-2 GeV, where high single-beam polarization has previously been observed. Further decisions regarding the proposal were deferred to the December meeting.
 

6. Discussion of the Proposal LEPC 94-9/P3 add.3: Proposal to Install Scintillating Tiles in the OPAL Endcap

The referee (B. Foster) presented his report on this proposal which aims to install a redundant forward trigger system in the OPAL endcap during the 95/96 shutdown. The committee recommended approval of the proposal subject to the availability of the necessary funds.
 

7. Discussion of the Memorandum LEPC 94-10/M109: Search for Cosmic Ray Coincidences with Detectors in the LEP tunnel

The committee recommended approval of the test experiment described in the present memorandum, which involves only the ALEPH detector and scintillator arrays placed nearby in the LEP tunnel.
 

8. Report from the LEP Co-ordinator

T. Camporesi reported on the LEP schedule for the remainder of 1994 and the 94/95 shutdown, and presented the proposed schedule for 1995. The second floating LEP technical stop is scheduled for 24 October - 4 November 1994. The dates of the 94/95 LEP shutdown are 5 December 1994 - 14 April 1995. Following cold check-out, LEP will start up on 24 April 1995.

Two long technical stops -each lasting about 3 weeks- are foreseen during 1995 for the installation of superconducting cavities for LEP 2. This would allow the possibility of high energy running by the end of the 1995 run. The proposed date for LEP shutdown is 20 November 1995. This shutdown starts about two weeks earlier than normal in order to accommodate the extensive repairs and installation required by L3 in preparation for LEP 2 operation. The totals are 118 days of physics in 1995 and 34 days of MD; the corresponding numbers for 1994 are 128 days and 37 days, respectively.
 

9. LEPC meetings in 1994 and 1995

The date of the next meeting of the LEPC will be Friday 9 December. The tentative dates of LEPC meetings at CERN in 1995 are:

31 January - 1 February,
6 - 7 June, and
14 - 15 November.
The Cogne meeting of the LEPC to review LEP physics and plans has been brought forward by one week to 17 - 23 September 1995 in order to avoid conflict with other CERN committee meetings.                                                                                                    J. Kirkby DOCUMENTS RECEIVED

[1] Report from the Bunch Train Working Group (T. Camporesi et al.), LEPC 94-13/M110.